Aeronautical

What is the maximum amount of quantization error that can occur in a 10bit system in the input analog voltage is from 0V to 6V?
a) 10mV
b) 6mV
c) 0.5mV
d) 1mV
Answer #1

Answer: b
Explanation: The minimum voltage step input is 6/(2N-1) = 6/1023 =5.865mV ≈6mV. This is the maximum error that can occur during the conversion process.

answered by: prof,amjad abdullah